OpenStudy (anonymous):

9x/x^2-4 when x = 1/2 The computer keeps getting -6/5 and I have NO idea how they are getting this answer.

OpenStudy (anonymous):

subtitute value of 'x=1/2' in equation ans= -6/5

OpenStudy (anonymous):

I did substitute but for some reason keep getting the wrong answer

OpenStudy (anonymous):

I'm getting 4.5/-3.75

jimthompson5910 (jim_thompson5910):

One method is to evaluate 4.5/-3.75 with a calculator to get 4.5/-3.75 = -1.2 Then convert that decimal into a fraction -1.2 = -(1 + 0.2) -1.2 = -(1 + 2/10) -1.2 = -(10/10 + 2/10) -1.2 = -(10+2)/10 -1.2 = -12/10 -1.2 = -6/5

OpenStudy (anonymous):

(9)(1/2) / (1/2)^2 -4 (9/2) / (1/4 - 4) (9/2) / (1-16 /4) 9 x 4 / 2 x (-15) cancellations: 9x2 / (-15) 18 / -15 cancellations: 6 / -5 or -6/5

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

now answer this: a sequence of numbers begin with the numbers -10, 20, 30,..., and each term afterward is the product of the preceding three terms. how many of the first 101 terms of this sequence are positive?

OpenStudy (anonymous):

plz